OpenStudy (ikou):

Please help Steve manages a skate park and he has been analyzing the attendance data. Steve finds that the number of visitors increases exponentially as the temperature increases, and this situation is represented by the function f(x) = 5x. Steve also finds a linear equation that models the number of people who leave the park early depending on the change in temperature, and it is represented by f(x) = −x + 6. The graph of the two functions is below. Find the solution to the two functions and explain what the solution represents.

OpenStudy (ikou):

The graph > https://gyazo.com/9363edf97d7607c2a157f3d93774dd31

OpenStudy (ikou):

Thanks for the help everyone

OpenStudy (ikou):

Outstanding

OpenStudy (gnorris):

Temperature is plotted on the x-axis and the number of visitors to the park and the number of visitors who leave the park early is plotted on the y-axis. Solve 5^x = -x+6 5^x + x - 6 = 0 x=1 is a solution by inspection. when x=1, y=5 when the temperature change is 1 degree, the number of people who leave the park early and the number of people who visit the park are equal.

OpenStudy (ikou):

I deem you the most helpful person on this website

OpenStudy (gnorris):

I like to help people well try and help as much as I can